OpenStudy (anonymous):

QUICK EASY QUESTION! WILL GIVE MEDAL! Simplify : (3^6)^18 a. 3^3 b. 3^12 c. 3^24 d. 3^108

OpenStudy (anonymous):

um try making it easier to read I might be able to help

TheSmartOne (thesmartone):

\[(a^x)^y=a ^{xy}\]

OpenStudy (igreen):

Use the exponent rule: \((x^y)^z = x^{y \cdot z}\)

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

\[(3^6)^{18}=3^{6\times18}\]

OpenStudy (igreen):

Applying that rule will give us: \((3^6)^{18} \rightarrow 3^{6 \cdot 18} \rightarrow 6^{108}\)

TheSmartOne (thesmartone):

So what is 6x18?

TheSmartOne (thesmartone):

@iGreen I thing you meant to write \[3^{108}\]

OpenStudy (anonymous):

ohhh so its 3^108?

TheSmartOne (thesmartone):

Yes.

OpenStudy (anonymous):

Thank u!

TheSmartOne (thesmartone):

Also note that \(x^y \times x^z=x ^{y+z}\)

OpenStudy (igreen):

Oh yeah my bad..xD